A wire of resistance R is elongated n-fold to make a new uniform wire. The resistance of new wire
(A) nR (B) n2
R (C) 2nR (D) 2n2
R

Grade:

2 Answers

Arun
3 years ago
Resistance of wire initially          R=Aρl            
Using  A=V/l        R=Vρl2           
New length wire after stretching    l=nl
New resistance      R=Vρ(nl)2=
Vikas TU
3 years ago
Wire of resistance R is elongated n fold to make a new uniform wire.So its radius is reduced by n times.
Let l  and r be the original length and radius of the original wire.So,
Original Resistance of the wire , R = ρl/πr^2
Now radius is reduced by n times ,so new radius is r' = r/n
New resistance,R' = ρl/πr'^2 = ρl/π(r/n)^2
= n^2ρl/πr^2 = n^2R